OpenStudy (anonymous):

each time sally goes for a run, she runs 2km more than david. Sally runs seven days a week and David runs six days a week. Sally runs 22km a week more than David. How far does each run per day?

OpenStudy (anonymous):

s: how much sally runs d: how much david runs 7s=6d+22 s=d+2 7(d+2)=6d+22 7d+14=6d+22 d=8 s=10